The longest side of a triangle is 16 less than the sum of the other two. The shortest side is three times the difference of the other two. The Perimeter is 150. Write the system of equation.
Answer by cleomenius(959)

Let
a = longest
b = middle
c = smallest
=========================================================
Set up our equations.
-a +b + c = 16
3a - 3b -c = 0
a + b + c = 150
=======================================================
multiply 1 * 3;add 1 to 2
-a +b + c = 16
2c = 48
a + b + c = 150
c = 24
=============================================
plug c into 1 and 3,
-a +b + 24= 16
2c = 48
a + b + 24= 150
multiply 1 by -1, add 1 to 3
-a +b + 24= 16
2c = 48
2a + 0= 134
a = 67 degrees.
========================================================
Solve for b.
67 + 24 = 91 ; 150 - 91 = 59
b = 59 degrees.
=======================================================
a = 67 degrees.
b = 59 degrees.
c = 24 degrees.
Now, to check the system.
59 + 24 = 83; 83 - 16 = 67.
67 - 59 = 8; 8*3 = 24, which is the smallest angle.
67 + 59 + 24 = 150.
This system does check.
Cleomenius.